March 7, 2023
Return to the calendar.
RC Meeting Information
Current RC: Iuliia Skorodumina.
No RC meeting today.
Monday Day (03/06) was a Maintenance Day. The repair of the 4th pass separator was being performed during the Day shift, while the separator phasing was scheduled for Swing (see PD update: https://logbooks.jlab.org/entry/4143986). The hall went to Controlled Access. The following work was being done in the hall.
Jerry and the team: pump SHMS transfer line.
Ellen replaced the relay for SHMS JT2 (https://logbooks.jlab.org/entry/4143930).
Jixie, Michael, and Jimmy: work on the solid detector.
Monday Swing (03/06). The repair of the separator was complete and the phasing commenced. By 9pm it was decided to pause the phasing for the night and to pick up on Tuesday (https://logbooks.jlab.org/entry/4144141). The beam was magnetically separated to Halls B, C, and D to run overnight. The data-taking was resumed at 30 uA first, and by 10:30pm we were getting 60 uA. Swing Shift summary: https://logbooks.jlab.org/entry/4144077 .
Tuesday Owl (03/07). Taking data overnight mostly at 60 uA with decent ABU (https://logbooks.jlab.org/entry/4144349). Fast Feedback system was turned on at 5:20 am (https://logbooks.jlab.org/entry/4144291).
For the current setting (Pm=800 MeV/c), we have collected a total of 14575 mC integrated charge out of the planned 15696 mC, which corresponds to 93%.
The 4th pass separator phasing will continue during the Day shift today and is scheduled to start at 9am, so the beam will be taken away.
Current Run Plan
We will need a few more runs to finish the Pm=800 MeV/c setting and then will move to the next setting with Pm=900 MeV/c (probably some time during Swing today depending on beam availability).